Please see the two pics, one is the question and other one is the formula, please show me the steps
currently A coupon bond that has a $1,000 par value and a coupon rate of 10%. The bond is selling for $1,250 and has eight years to maturity. What is the bonds yield to maturity? (An equation is sufficient.) 3,
3. Coupon Bond Using the same strategy used for the fixed-payment loan P price of coupon bond C yearly coupon payment F face value of the bond n years to maturity date 1ti (1+i) (1+i)
Answer #1

Maturity = 8 Years (n)
Price = 1250(P)
Par Value = 1000(F)
Coupon = Coupon Rate* Par Value = 10%*1000 = 100(C)

P = C/(1+i) + C/(1+i)2 + C/(1+i)3 + ....... + 1000/(1+i)10
1250 = 100/(1+i) + 100/(1+i)2 + 100/(1+i)3 ......+ 1000/(1+i)10
1250 = 1000(1-(1+i)-10)/i + 1000/(1+i)10 (Sum of Coupons using PV of annuity formula)

Hit and Trial Method
Applying i =9% , We get RHS value = 1064.18
i = 8% , We Get RHS value = 1134.20
i = 7% , We Get RHS value = 1210.71
i= 6% We get RHS = 1294.50
I = 6.5% We get RHS = 1251.61
AT i = 6.52% We get RHS = 1250

Hence YTM = 6.52%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well